An elastic cord is 61 cm long when a weight of 75 N hangs from it but is 85 cm long when a weight of 210 N hangs from it. What is the “spring” constant k of this elastic cord?


1
Expert's answer
2021-03-16T11:37:54-0400

Let the length of the elastic cord at rest be "x" meters. Then, according to the Hooke's Law we can write the spring constant "k" for both cases:


"k=\dfrac{F_1}{\Delta x_1}=\dfrac{75}{0.61-x},""k=\dfrac{F_2}{\Delta x_2}=\dfrac{210}{0.85-x}."

Then, we get:


"\dfrac{75}{0.61-x}=\dfrac{210}{0.85-x},""75(0.85-x)=210(0.61-x),""135x=64.35,""x=\dfrac{64.35\ N\cdot m}{135\ N}=0.48\ m."

Therefore, we can calculate the spring constant "k":


"k=\dfrac{F_1}{\Delta x_1}=\dfrac{75\ N}{0.61\ m-0.48\ m}=577\ \dfrac{N}{m}."